Please help me- I do not understand how to solve for k.
(-9/2)k + (-6/5) = 9 - (9/8)k
Thank you!

Answer by jim_thompson5910(35256)

(-9/2)k + (-6/5) = 9 - (9/8)k

40*(-9/2)k + 40*(-6/5) = 40*9 - 40*(9/8)k ... Multiply EVERY term by the LCD 40 to clear out the fractions.

20*(-9)k + 8*(-6) = 40*9 - 5*(9)k

-180k - 48 = 360 - 45k

-180k=360-45k+48

-180k+45k=360+48

-135k=360+48

-135k=408

k=(408)/(-135)

k=-136/45

So the solution is k = -136/45